OpenStudy (anonymous):

y varies indirectly as twice x. When x = 1, y = 3. Find y when x = 4.

jimthompson5910 (jim_thompson5910):

"y varies indirectly as twice x" means that \[\Large y = \frac{k}{2x}\] Now plug in the given info to find k \[\Large y = \frac{k}{2x}\] \[\Large 3 = \frac{k}{2(1)}\] \[\Large 3 = \frac{k}{2}\] \[\Large 3*2 = k\] \[\Large 6 = k\] \[\Large k = 6\] So the equation is \[\Large y = \frac{6}{2x}=\frac{3}{x}\], which is simply \[\Large y = \frac{3}{x}\] Now plug in x = 4 \[\Large y = \frac{3}{x}\] \[\Large y = \frac{3}{4}\] So when x = 4, y = 3/4